OpenStudy (anonymous):

simplify the expression x(sqrt)8x3 + x(sqrt)2 + 4(sqrt)2x

OpenStudy (anonymous):

\[x \sqrt{8 \times3} + 2\sqrt{2} + 4\sqrt{2x}\] ???

OpenStudy (anonymous):

is that what you are asking?

OpenStudy (anonymous):

except that 2 is supposed to be an x

OpenStudy (anonymous):

2xsqrt6 +xsqrt2 +4sqrt2x